Show steps on how to you get the answers.

Required information [The following information applies to the questions displayed below.) The equity sections for Atticus GrThe following transactions and events affected its equity during the year. Jan. 5 Declared a $0.50 per share cash dividend, dRequired: 1. How many common shares are outstanding on each cash dividend date? Jan. 5 Apr. 5 July 5 Oct. 5 Outstanding commo5. How much net income did the company earn this year? Net income
Answer #1

1) Outstanding shares

Jan 5 Apr 5 Jul 5 Oct 5
Outstanding share 35000 35000-3000 = 32000 32000 32000*1.20 = 38400

2) Cash dividend

Jan 5 Apr 5 Jul 5 Oct 5
Cash dividend 35000*.50 = 17500 32000*.50 = 16000 16000 38400*.50 = 19200

5) Net income = 360000-68700-64000-400000 = $172700